# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v2

@@ 4.2 SEIR
This second example introduces an exposed compartment for diseases with delayed onset, a version of the model was laid out by Cooke in 1966 [20]. The additional compartment $E$ represents exposed, usually meaning a person is infected with a disease but can’t infect others yet. The parameters $\Pi,\beta,\mu,\eta,$ and $\alpha$ represent the birth/immigration rate, the infection rate, the death/emigration rate, the onset of infectiousness/symptoms rate, and the recovery rate, respectively.
$\frac{dS}{dt}$ $=$ $\Pi-\beta SI-\mu S$ (8)
$\frac{dE}{dt}$ $=$ $\beta SI-\eta E-\mu E$ (9)
$\frac{dI}{dt}$ $=$ $\eta E-\alpha I-\mu I$ (10)
$\frac{dR}{dt}$ $=$ $\alpha I-\mu R$ (11)
The Equations 8-11 are mapped to a VAPN (Figure 3(a)) and stochastic Petri net (Figure 3(b)).
Figure 3: Two Petri net implementations of the SEIR ODE model given in Equations 8-11. Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
The first step in finding the $R_{0}$ of the PN systems in Figure 3 is to obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E^{*},I^{*},R^{*})=(\frac{\Pi}{\mu},0,0,0,0)$. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments, outlined in red in Figure 3 and as shown in Equation 2. Note that $i$ is with respect to the corresponding infected compartment and the corresponding row in the matrix, which we number. For this example, the $E$ compartment is 1 and the $I$ compartment is 2. The terms for the VAPN in Figure 3(a) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=0$. Similarly the terms for the SPN in Figure 3(b) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=\beta SI-\beta SI=0$. Thus,
@@
for VAPN as shown in Figure 3(a), and
@@
for the stochastic PN as shown in Figure 3(b).
From this we find the rate of change with respect to $E$ in the first column and $I$ in the second column, and evaluate at the DFE. This gives the same result for both Petri net options:
@@
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another, as outlined in green in Figure 3. Then for $\mathcal{V}_{1}$ the terms $w(I_{1},t_{j,...,k})=\eta E+\mu E$, $w(t_{N_{m},...,N_{l}}=0,I_{1})$, and $-w(t_{I_{2}},I_{1})=0$. For $\mathcal{V}_{2}$ the terms $-w(t_{I_{1}},I_{2})=-\eta E$, $w(I_{2},t_{j,...,k})=\alpha I+\mu I$, and $w(t_{N_{m},...,N_{l}},I_{2})=0$. Both $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are the same for the VAPN and SPN in Figures 3(a) and 3(b) respectively. Thus,
@@
We then find the rate of change with respect to the infected place $E$ in the first column and $I$ in the second column, and we evaluate it at the DFE. From this we find the inverse to give
$V$ $=$ $\begin{bmatrix}\eta+\mu&0\\ -\eta&\alpha+\mu\end{bmatrix}$
$V^{-1}$ $=$ $\begin{bmatrix}\frac{1}{\eta+\mu}&0\\ \frac{a}{(n+u)(a+u)}&\frac{1}{a+\mu}\end{bmatrix}.$
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta\Pi\eta}{\mu(\alpha+\mu)(\eta+\mu)}$
This is analytically the same $R_{0}$ expression that can be derived for the ODE using the next-generation matrix method.
@@ 4.3 SEEIR
This third example introduces parallel paths of infected compartments. The model and subsequent $R_{0}$ value was outlined by Diekmann [5]. The parameters $\mu,\beta,p,v_{1},v_{2},$ and $\gamma$ represent birth/death rate, transmission rate, rate of exposure to infection, rate of becoming infectious for $E_{1}$, rate of becoming infectious for $E_{2}$, and rate of recovery, respectively.
$\frac{dS}{dt}$ $=$ $\mu N-\beta\frac{SI}{N}-\mu S$ (13)
$\frac{dE_{1}}{dt}$ $=$ $p\beta\frac{SI}{N}-v_{1}E_{1}-\mu E_{1}$ (14)
$\frac{dE_{2}}{dt}$ $=$ $(1-p)\beta\frac{SI}{N}-v_{2}E_{2}-\mu E_{2}$ (15)
$\frac{dI}{dt}$ $=$ $v_{1}E_{1}+v_{2}E_{2}-\gamma I-\mu I$ (16)
$\frac{dR}{dt}$ $=$ $\gamma I-\mu R$ (17)
The Equations 13-17 are mapped to a variable arc weight Petri net to yield Figure 4. We will analyze only the Variable Arc Weight Petri net (VAPN) implementation going forward for brevity and since it is the more novel approach.
Figure 4: This SEEIR Petri net model is mapped directly from Equations 13-17. The arc weights used in forming matrix $F_{i}(x)$ are outlined in red and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
The first step in finding the $R_{0}$ of the PN system in Figure 4 is to obtain the DFE by setting the token levels of the infected places to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E_{1}^{*},E_{2}^{*},I^{*},R^{*})=(S^{*},0,0,0,0)$. Note since $N=S+E_{1}+E_{2}+I+R$, $S^{*}=N$, where $N$ is the total population. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments of $E_{1},E_{2},$ and $I$ as shown in Equation 2. Note that $i$ refers to the corresponding infected compartment and the corresponding row in the matrix, which we have numbered. For this example, the $E_{1}$ compartment is 1, the $E_{2}$ compartment is 2 and the $I$ compartment is 3.
@@
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another as shown in Equation 3. This yields
@@
$V$ $=$ $\begin{bmatrix}v_{1}+\mu&0&0\\ 0&v_{2}+\mu&0\\ -v_{1}&-v_{2}&\gamma+\mu\end{bmatrix}$
@@
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue.
@@
$R_{0}$ $=$ $\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\left(\frac{pv_{1}}{v_{1}+\mu}+\frac{(1-p)v_{2}}{v_{2}+\mu}\right)\frac{\beta}{\gamma+\mu}$
This is analytically the same $R_{0}$ expression found in Equation 2.10 of [5].
4.4 Basic COVID Model
This next example was laid out as a basic COVID-19 model to help with forecasting once the appropriate compartments for the disease had been classified [21]. This example introduces more complex equations for $F$, more interactions involved in $V$, and application to a specific disease. By applying our method to a compartmental COVID-19 model, we demonstrate its practical utility in analyzing contemporary and complex epidemiological challenges, where multiple infectious states with varying transmission dynamics play a crucial role. The compartments $S,E,I_{a},I_{s},I_{h},$ and $R$ represent susceptible, exposed, infected asymptomatic, infected symptomatic, infected hospitalized, and recovered, respectively. The parameters $\beta_{a},\beta_{S},$ and $\beta_{h}$ represent the rate of infection due to asymptomatic, symptomatic, and hospitalized individuals, respectively. The parameters $\sigma,r,\gamma_{a},\gamma_{s},\varphi_{S},\delta_{s},\gamma_{h},$ and $\delta_{h}$ represent rate of becoming infectious, the percentage of people who are asymptomatic, rate of recovery for asymptomatic infected, rate of recovery for symptomatic infected, rate of hospitalization of symptomatic, death rate of symptomatic, rate of recovery for hospitalized and death rate of hospitalized, respectively.
$\frac{dS}{dt}$ $=$ $-\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S$ (18)
$\frac{dE}{dt}$ $=$ $\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S-\sigma E$ (19)
$\frac{dI_{a}}{dt}$ $=$ $r\sigma E-\gamma_{a}I_{a}$ (20)
$\frac{dI_{S}}{dt}$ $=$ $(1-r)\sigma E-(\varphi_{s}+\gamma_{s}+\delta_{s})I_{s}$ (21)
$\frac{dI_{h}}{dt}$ $=$ $\varphi_{s}I_{s}-(\gamma_{h}+\delta_{h})I_{h}$ (22)
$\frac{dR}{dt}$ $=$ $\gamma_{a}I_{a}+\gamma_{s}I_{s}+\gamma_{h}I_{h}$ (23)
The Equations 18-23 are mapped to a variable arc weight Petri net to yield Figure 5.
@@ 4.4 Basic COVID Model
Figure 5: A variable arc weight COVID Petri net mapped from Equations 18-23. The arc weights used in forming matrix $F_{i}(x)$ are outlined in red and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
From Figure 5, we can see that there are four infected places ($E,I_{a},I_{S},I_{h}$) which are the first, second, third, and fourth infected compartments, respectively. Our resulting $F$ and $V$ matrices will therefore be four by four. The disease free equilibrium is given by $(S^{*},E^{*},I_{a}^{*},I_{s}^{*},I_{h}^{*},R^{*})=(S^{*},0,0,0,R^{*})$. We defined $N=S+E+I_{a}+I_{S}+I_{h}+R$, so $N^{*}=S^{*}+R^{*}$. By inspecting the PN model we see that the term $\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S$ is the only way the Susceptible place leads to an infected place. The transmission matrix starts with the resulting $\mathcal{F}$ using Equation 2 is
$\mathcal{F}_{1,2,3,4}(x)$ $=$ $\begin{bmatrix}\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$
Taking the partial derivative with respect $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
@@
$F$ $=$ $\begin{bmatrix}0&\frac{\beta_{a}S^{*}}{N^{*}}&\frac{\beta_{S}S^{*}}{N^{*}}&\frac{\beta_{h}S^{*}}{N^{*}}\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$

### v3

@@ 4.2 SEIR
This second example introduces an exposed compartment for diseases with delayed onset. Cooke laid out a version of the model Cooke in 1966 Cooke (1966). The additional compartment $E$ represents exposed individuals, usually meaning a person is infected with a disease but cannot infect others yet. The parameters $\Pi,\beta,\mu,\eta,$ and $\alpha$ represent the birth/immigration rate, the infection rate, the death/emigration rate, the onset of infectiousness/symptoms rate, and the recovery rate, respectively. The SEIR system is described by the following system of equations:
$\frac{dS}{dt}$ $=\Pi-\beta SI-\mu S,$ (7)
$\frac{dE}{dt}$ $=\beta SI-\eta E-\mu E,$ (8)
$\frac{dI}{dt}$ $=\eta E-\alpha I-\mu I,$ (9)
$\frac{dR}{dt}$ $=\alpha I-\mu R.$ (10)
The Equations (7)–(10) are mapped to a VAPN (Figure 3a) and stochastic Petri Net (Figure 3b).
Figure 3: Two Petri Net implementations of the SEIR ODE model given in Equations (7)–(10). Option (a) uses a discrete-time, variable arc weight PN. Option (b) uses a continuous-time, stochastic Markov Chain PN. The arcs and arc weights used in forming matrix $F_{i}(x)$ are outlined and highlighted in red, and the arcs and arc weights used in forming matrix $V_{i}(x)$ are outlined and highlighted in green.
To find the $R_{0}$ of the PN systems in Figure 3 we follow the same procedure outlined in Section 4.1, we first obtain the DFE. By setting the infected tokens to zero and solving for the non-infected steady state, we obtain $(S^{*},E^{*},I^{*},R^{*})=(\frac{\Pi}{\mu},0,0,0,0)$. Then, the matrix $F_{i}(x)$ can be found by examining the arc weights of arcs entering the infected compartments, outlined in red in Figure 3, as shown in Equation (2). Note that $i$ refers to the corresponding infected compartment and the corresponding row in the matrix, which we have numbered. For this example, the $E$ compartment is 1 and the $I$ compartment is 2. The terms for the VAPN in Figure 3a within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,\ldots,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,\ldots,k}},I_{1})=0$. Similarly the terms for the SPN in Figure 3b within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,\ldots,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,\ldots,k}},I_{1})=\beta SI-\beta SI=0$. Thus,
@@
for VAPN as shown in Figure 3a, and
@@
for the stochastic PN as shown in Figure 3b.
As $E$ is our first infected compartment and $I$ is the second, we find the rate of change with respect to $E$ in the first column and $I$ in the second column, and evaluate at the DFE. This gives the same result for both Petri Net options:
@@
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another, as outlined in green in Figure 3. Then for $\mathcal{V}_{1}$ the terms $w(I_{1},t_{j,\ldots,k})=\eta E+\mu E$, $w(t_{N_{m},\ldots,N_{l}}=0,I_{1})$, and $w(t_{I_{2}},I_{1})=0$. For $\mathcal{V}_{2}$ the terms $w(t_{I_{1}},I_{2})=-\eta E$, $w(I_{2},t_{j,\ldots,k})=\alpha I+\mu I$, and $w(t_{N_{m},\ldots,N_{l}},I_{2})=0$. Both $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are the same for the VAPN and SPN in Figures 3a and 3b respectively. Thus,
@@
Next, we find the rate of change with respect to the infected place $E$ in the first column and $I$ in the second column, and we evaluate it at the DFE. From this, we find the inverseto give
$V$ $=\begin{bmatrix}\eta+\mu&0\\ -\eta&\alpha+\mu\end{bmatrix}$,
$V^{-1}$ $=\begin{bmatrix}\frac{1}{\eta+\mu}&0\\ \frac{\alpha}{(\eta+\mu)(\alpha+\mu)}&\frac{1}{\alpha+\mu}\end{bmatrix}$.
Now that we have the $F$ and $V$ matrices, we can find $R_{0}$ by multiplying them and then determining the dominant eigenvalue. Thus,
$R_{0}$ $=\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\frac{\beta\Pi\eta}{\mu(\alpha+\mu)(\eta+\mu)}.$
This is analytically the same $R_{0}$ expression that can be derived for the ODE using the next-generation matrix method. This demonstrates a key strength of our framework, which is its ability to extract the fundamental epidemiological dynamics relevant to $R_{0}$ irrespective of the specific Petri Net formalism used for implementation, whether deterministic VAPNs or stochastic SPNs.
@@ 4.3 SEEIR
This third example introduces parallel paths of infected compartments. The model and subsequent $R_{0}$ value were outlined by Diekmann Diekmann et al. (2010). The parameters $\mu,\beta,p,v_{1},v_{2},$ and $\gamma$ represent birth/death rate, transmission rate, rate of exposure to infection, rate of becoming infectious for $E_{1}$, rate of becoming infectious for $E_{2}$, and rate of recovery, respectively. The SEEIR system is described by the following system of equations:
$\frac{dS}{dt}$ $=\mu N-\beta\frac{SI}{N}-\mu S,$ (11)
$\frac{dE_{1}}{dt}$ $=p\beta\frac{SI}{N}-v_{1}E_{1}-\mu E_{1},$ (12)
$\frac{dE_{2}}{dt}$ $=(1-p)\beta\frac{SI}{N}-v_{2}E_{2}-\mu E_{2},$ (13)
$\frac{dI}{dt}$ $=v_{1}E_{1}+v_{2}E_{2}-\gamma I-\mu I,$ (14)
$\frac{dR}{dt}$ $=\gamma I-\mu R.$ (15)
The Equations (11)–(15) are mapped to a variable arc weight Petri Net to yield Figure 4. We will analyze only the Variable Arc Weight Petri Net (VAPN) implementation going forward, for brevity, as it is the more novel approach. The corresponding SPN models would be analyzed using the same principles, relying on the net sum of continuous-time rate equations for each transition, as demonstrated for the SEIR model, yielding identical final $F$ and $V$ matrices at the DFE.
Figure 4: This SEEIR Petri Net model is mapped directly from Equations (11)–(15). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
The first step in finding the $R_{0}$ of the PN system in Figure 4 is to obtain the DFE by setting the token levels of the infected places to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E_{1}^{*},E_{2}^{*},I^{*},R^{*})=(S^{*},0,0,0,0)$. Note that since $E_{1}+E_{2}+I+R, S=NS^{*}=N$, where $N$ is the total population. Then, the matrix $F_{i}(x)$ can be found by examining the arc weights of arcs entering the infected compartments of $E_{1},E_{2},$ and $I$, as shown in Equation (2). Note that $i$ refers to the corresponding infected compartment and the corresponding row in the matrix, which we have numbered. For this example, the $E_{1}$ compartment is 1, the $E_{2}$ compartment is 2, and the $I$ compartment is 3.
@@
For $\mathcal{V}$, we analyze the Petri net to determine arcs leaving infected places as shown in Equation (3). This yields
@@
$V$ $=\begin{bmatrix}v_{1}+\mu&0&0\\ 0&v_{2}+\mu&0\\ -v_{1}&-v_{2}&\gamma+\mu\end{bmatrix}$,
@@
Now that we have $F$ and $V$ matrices, we can find $R_{0}$ by multiplying and finding the dominant eigenvalue. Thus
@@
$R_{0}$ $=\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\left(\frac{pv_{1}}{v_{1}+\mu}+\frac{(1-p)v_{2}}{v_{2}+\mu}\right)\frac{\beta}{\gamma+\mu}.$
This is analytically the same $R_{0}$ expression found in Equation 2.10 of Diekmann et al. (2010).
4.4 SVEIR
While $R_{0}$ is the average number of secondary infections produced by a single infectious individual introduced into a completely susceptible population Diekmann et al. (2010), there is also the notion of effective reproduction number of $R_{e}$ which is the average number of secondary infections produced by a single infectious individual at a specific point in time ($t$) Cori et al. (2013). With this in mind, we move to an extension of the SEIR model by Bugalia et al. with a compartment for Vaccinated individuals $V$ and $N$ is the sum of all compartments Bugalia et al. (2023). With the compartment $V$ we can look at the effective reproduction number when vaccination and non-pharmaceutical interventions are occurring. The parameters $\beta_{t},\xi_{t},\gamma,$ and $\varphi$ represent the contact rate of individuals (time dependent), vaccination rate (time dependent), the disease re-susceptibility rate, and the vaccine re-susceptibility rate, respectively. The parameters $\alpha,$ and $\delta$ represent the recovery rate of infected individuals and disease death rate, respectively. The vaccine efficacy is described as $1-\sigma$ where vaccinated people are infected at the rate $\sigma\beta_{t}$. The SVEIR system is described by the following system of equations:
$\frac{dS}{dt}$ $=-\frac{S}{N}\beta_{t}I+\gamma R-\xi_{t}S+\varphi V$ (16)
$\frac{dV}{dt}$ $=-\frac{V}{N}\sigma\beta_{t}I+\xi_{t}S-\varphi V$ (17)
$\frac{dE}{dt}$ $=\beta_{t}I(\frac{S}{N}+\sigma\frac{V}{N})-\alpha E$ (18)
$\frac{dI}{dt}$ $=\alpha E-(\alpha+\delta)I$ (19)
$\frac{dR}{dt}$ $=\alpha E-\gamma R$ (20)
The Equations (17)–(20) are mapped to a VAPN in Figure 5.
@@ 4.4 SVEIR
Figure 5: A variable arc weight SVEIR Petri Net mapped from Equations (17)–(20). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
The steps for finding the effective reproduction number $R_{e}$ of the PN systems remains the same as the reproduction number for ODEs if Lemma 1 from van den Driessche and Watmough Van den Driessche and Watmough (2002) holds true. If the assumptions we laid out for Petri Nets (A1–A5) also hold true for our Petri Net model, we then find the effective reproduction number for Petri Nets in the same method as we would for the reproduction number. Thus, we first obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},V^{*},E^{*},I^{*},R^{*})=(N\frac{\varphi}{\varphi+\xi_{t}},N\frac{\xi_{t}}{\varphi+\xi_{t}},0,0,0)$. Note that since $N=S+V+E+I+R$, $S^{*}+V^{*}=N$, where $N$ is the total population. Then, the matrix $F_{i}(x)$ can be found by examining the arc weights of arcs entering the infected compartments of $E$ and $I$, as shown in Equation (2). Note that $i$ refers to the corresponding infected compartment and the corresponding row in the matrix, which we have numbered. For this example, the $E$ compartment is 1 and the $I$ compartment is 2. Thus,
$\mathcal{F}_{1,2}(x)$ $=\begin{bmatrix}\beta_{t}I(\frac{S}{N}+\sigma\frac{V}{N})&\beta_{t}I(\frac{S}{N}+\sigma\frac{V}{N})\\ 0&0\end{bmatrix}$,
for VAPN as shown in Figure 5.
We then find the rate of change with respect to $E$ in the first column and $I$ in the second column, and evaluate at the DFE, yielding
$F$ $=\begin{bmatrix}0&\beta_{t}(\frac{\varphi+\sigma\xi_{t}}{\varphi+\xi_{t}})\\ 0&0\\ \end{bmatrix}$.
For $\mathcal{V}$, we find arcs coming out of infected places, outlined in green to help. This yields
$\mathcal{V}_{1,2}(x)$ $=\begin{bmatrix}\alpha E&0\\ -\alpha E&(\alpha+\delta)I\end{bmatrix}$.
$V$ $=\begin{bmatrix}\alpha&0\\ -\alpha&\alpha+\delta\end{bmatrix}$,
$V^{-1}$ $=\begin{bmatrix}\frac{1}{\alpha}&0\\ \frac{1}{\alpha+\delta}&\frac{1}{\alpha+\delta}\end{bmatrix}$.
With the $F$ and $V$ matrices in hand, we can find $R_{0}$. Thus,
@@
$R_{0}$ $=\varphi\begin{bmatrix}FV^{-1}\end{bmatrix}=\varphi\begin{bmatrix}\frac{\beta_{t}}{\alpha+\delta}(\frac{\varphi+\sigma\xi_{t}}{\varphi+\xi_{t}})&\frac{\beta_{t}}{\alpha+\delta}(\frac{\varphi+\sigma\xi_{t}}{\varphi+\xi_{t}})\\ -\alpha&\alpha+\delta\end{bmatrix}=\frac{\beta_{t}}{\alpha+\delta}(\frac{\varphi+\sigma\xi_{t}}{\varphi+\xi_{t}})$
This is analytically the same $R_{e}$ expression found in Bugalia et al. Bugalia et al. (2023). Note that as the vaccine efficacy increases, vaccine rate increases, or contact rate decreases, the effective reproduction number decreases.
4.5 Basic COVID-19 Model
The next example was laid out as a basic COVID-19 model to aid in forecasting once the appropriate compartments for the disease had been identified Gumel et al. (2021). This example introduces more complex equations for $F$, involving additional interactions in $V$, and applies it to a specific disease. By applying our method to a compartmental COVID-19 model, we demonstrate its practical utility in analyzing contemporary and complex epidemiological challenges, where multiple infectious states with varying transmission dynamics play a crucial role. The compartments $S,E,I_{a},I_{s},I_{h},$ and $R$ represent susceptible, exposed, infected asymptomatic, infected symptomatic, infected hospitalized, and recovered, respectively. The parameters $\beta_{a},\beta_{S},$ and $\beta_{h}$ represent the rates of infection due to asymptomatic, symptomatic, and hospitalized individuals, respectively. The parameters $\sigma,r,\gamma_{a},\gamma_{s},\varphi_{S},\delta_{s},\gamma_{h},$ and $\delta_{h}$ represent rate of becoming infectious, the percentage of asymptomatic people, rate of recovery for asymptomatic infected, rate of recovery for symptomatic infected, rate of hospitalization of symptomatic, death rate of symptomatic, rate of recovery for hospitalized and death rate of hospitalized, respectively. The Basic COVID-19 system is described by the following system of equations:
$\frac{dS}{dt}$ $=-\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S,$ (21)
$\frac{dE}{dt}$ $=\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S-\sigma E,$ (22)
$\frac{dI_{a}}{dt}$ $=r\sigma E-\gamma_{a}I_{a},$ (23)
$\frac{dI_{S}}{dt}$ $=(1-r)\sigma E-(\varphi_{s}+\gamma_{s}+\delta_{s})I_{s},$ (24)
$\frac{dI_{h}}{dt}$ $=\varphi_{s}I_{s}-(\gamma_{h}+\delta_{h})I_{h},$ (25)
$\frac{dR}{dt}$ $=\gamma_{a}I_{a}+\gamma_{s}I_{s}+\gamma_{h}I_{h}.$ (26)
The Equations (21)–(26) are mapped to a variable arc weight Petri Net to yield Figure 6.
Figure 6: A variable arc weight COVID-19 Petri Net mapped from Equations (21)–(26). The arc weights used in forming matrix $F_{i}(x)$ are outlined in red, and the arc weights used in forming matrix $V_{i}(x)$ are outlined in green.
From Figure 6, we can see that there are four infected places ($E,I_{a},I_{S},I_{h}$) which are the first, second, third, and fourth infected compartments, respectively. Our resulting $F$ and $V$ matrices will therefore be four by four. The disease free equilibrium is given by $(S^{*},E^{*},I_{a}^{*},I_{s}^{*},I_{h}^{*},R^{*})=(S^{*},0,0,0,R^{*})$. We defined $N=S+E+I_{a}+I_{S}+I_{h}+R$, so $N^{*}=S^{*}+R^{*}$. By inspecting the PN model, we see that the term $\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S$ is the only way the Susceptible place leads to an infected place. The transmission matrix starts with the resulting $\mathcal{F}$ using Equation (2) is
$\mathcal{F}_{1,2,3,4}(x)$ $=\begin{bmatrix}\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S&\frac{\beta_{a}I_{a}+\beta_{S}I_{S}+\beta_{h}I_{h}}{N}S\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$.
Taking the partial derivative with respect to $E,I_{a},I_{S},I_{h}$ for the first, second, third, and fourth columns, respectively, gives us
$F$ $=\begin{bmatrix}0&\frac{\beta_{a}S^{*}}{N^{*}}&\frac{\beta_{S}S^{*}}{N^{*}}&\frac{\beta_{h}S^{*}}{N^{*}}\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$,